OpenStudy (anonymous):

Given f(x) = x2 + 6x and g(x) = 3x2 + 12x, find (f – g)(x).

OpenStudy (anonymous):

here is a simple breakdown: f(x)=(x^2 +6x) g(x)=(3x^2 +12x) (f-g)(x)= (f(x))-(g(x)) (f-g)(x)= (x^2 +6x)-(3x^2 +12x) can you get it from there?

OpenStudy (anonymous):

combine like terms right?

OpenStudy (johnweldon1993):

Correct...

OpenStudy (anonymous):

4x^2+18x

OpenStudy (anonymous):

oops wait

OpenStudy (anonymous):

2x^2+6x?

OpenStudy (johnweldon1993):

Remember first you need to distribute that '-' sign...

OpenStudy (anonymous):

oh so then it's 4x^2+13x the second part?

OpenStudy (johnweldon1993):

so you'll have x² - 3x² = ...? and 6x - 12x = ...?

OpenStudy (anonymous):

Just as john said, that negative sign gets distributed, then it is just a matter of combining like terms.

OpenStudy (johnweldon1993):

So you have (x² +6x)-(3x² +12x) After distributing that '-' sign into the second parenthesis you have x² + 6x - 3x² - 12x Now combine those...

OpenStudy (anonymous):

ohh okay 2x^2-6x then

OpenStudy (johnweldon1993):

missing 1 thing still... x² - 3x² = ...?

OpenStudy (anonymous):

-2x^2 I mean

OpenStudy (johnweldon1993):

There we go...correct -2x² - 6x

OpenStudy (anonymous):

thanks!

OpenStudy (johnweldon1993):

No problem!